OpenStudy (bossimbacon):

Solve the formula for the variable B. V = (Bh)/3 explain your answer pls

OpenStudy (bossimbacon):

@nincompoop

OpenStudy (jadedry):

\[V = (Bh)/3\] Multiply V by 3 to cancel out the 3 on the right. \[3V = Bh\] Divide by h to isolate B. \[3V/h = B\]

OpenStudy (daggiemib):

3V = bh then divide each side by h to get: (3V)/h = b